### v3

@@ 2. Definitions and model setup
It is also shown that the mean absolute error of the $NI$-error associated with the empirical measure converges to 0 (uniformly in time), that if the solution of (2.4) is to converge in time, then the limit can be characterised as a certain fixed point, and is an $\varepsilon$-MNE for $L$ given that $\beta$ is above a specified threshold.
@@
(3.5) $I(\theta)=\frac{\beta}{4}\int_{0}^{T}\sup_{f:\langle f,\theta(t)\rangle\neq 0}\frac{\langle\dot{\theta}(t)-A_{\theta}^{*}\theta(t),f\rangle^{2}}{\langle|\nabla_{z}f|^{2},\theta(t)\rangle}dt,$
@@
It remains to move from the rate function (3.3) to the calculus of variations form (3.5). Starting with the control-formulation, we can use the contraction principle to obtain an LDP for $\{\gamma^{n}\}$ on $C([0,T]:\mathcal{P}(\mathcal{Z}))$. Let $I$ denote the corresponding rate function. Standard calculations suggest that it can be re-written as (3.5). With $A^{u}_{\theta}$ the generator associated with (3.4), and $(A^{u}_{\theta})^{*}$ the corresponding (formal) adjoint—with $A^{*}_{\theta}$ corresponding to the case $u\equiv 0$—the PDE characterisation of the dynamics (3.4) is
@@
See, e.g., [23, 16] for a heuristic description of how to go from (3) and the control formulation (3.3) to this version of the rate function. In the recent works [11, 12] the authors give, to the best of our knowledge, the first rigorous proof of this type of equivalence between the different types of rate functions, for both moderate and large deviations, in the setting of multi-scale processes. The above expression for $I$ is precisely (3.5), the prescribed form of the rate function. ∎
@@
The sequence $\{NI_{T}(\gamma^{n})\}_{n}$ converges almost surely in $C([0,T]:\mathbb{R})$ to $NI_{T}(\gamma)$, where $\gamma$ is the same as in Corollary 3.3.
@@
a.s., where $\tilde{C}$ is an upper bound on the $BL$-norm of $x\mapsto\int l(x,y)d\mu_{y}(y)$ (for any $\mu_{y}$) and $y\mapsto\int l(x,y)d\mu_{x}(x)$ (for any $\mu_{x}$). By the convergence $\gamma^{n}\to\gamma$, for a given $\tilde{\varepsilon}>0$, there is a $N_{\tilde{\varepsilon}}\in\mathbb{N}$ such that $d_{BL}(\gamma^{n},\gamma)\leq\tilde{\varepsilon}$ a.s. for all $n\geq N_{\tilde{\varepsilon}}$. This also implies the same property for the $t$-marginals: a.s., for any $t\in[0,T]$, $d_{BL}(\gamma^{n}_{t},\gamma_{t})\leq\tilde{\varepsilon}$ if $n\geq N_{\tilde{\varepsilon}}$. Therefore, since the constant $\tilde{C}$ does not depend on $n$ or $t$,

### v4

@@ 2. Definitions and model setup
It is also shown that the mean absolute error of the NINI-error associated with the empirical measure converges to 0 (uniformly in time), that if the solution of (2.4) is to converge in time, then the limit can be characterised as a certain fixed point, and is an $\varepsilon$-MNE for $L$ given that $\beta$ is above a specified threshold.
@@
(3.5) $I(\theta)=\frac{\beta}{4}\int_{0}^{T}\sup_{f:\langle|\nabla_{z}f|^{2},\theta(t)\rangle\neq 0}\frac{\langle\dot{\theta}(t)-A_{\theta}^{*}\theta(t),f\rangle^{2}}{\langle|\nabla_{z}f|^{2},\theta(t)\rangle}dt,$
@@
It remains to move from the rate function (3.3) to the calculus of variations form (3.5). Starting with the control-formulation, we can use the contraction principle to obtain an LDP for $\{\gamma^{n}\}$ on $C([0,T]:\mathcal{P}(\mathcal{Z}))$. Let II denote the corresponding rate function. Standard calculations suggest that it can be re-written as (3.5). With $A^{u}_{\theta}$ the generator associated with (3.4), and $(A^{u}_{\theta})^{*}$ the corresponding (formal) adjoint—with $A^{*}_{\theta}$ corresponding to the case $u\equiv 0$—the PDE characterisation of the dynamics (3.4) is
@@
See, e.g., [23, 16] for a heuristic description of how to go from (3) and the control formulation (3.3) to this version of the rate function. In the recent works [11, 12] the authors give, to the best of our knowledge, the first rigorous proof of this type of equivalence between the different types of rate functions, for both moderate and large deviations, in the setting of multi-scale processes. The above expression for II is precisely (3.5), the prescribed form of the rate function. ∎
@@
The sequence $\{NI_{T}(\gamma^{n})\}_{n}$ converges almost surely in $C([0,T]:\mathbb{R})$ to $NI_{T}(\gamma)$, where $\gamma$ is the same as in Corollary $3.$
@@
a.s., where $\tilde{C}$ is an upper bound on the BLBL-norm of $x\mapsto\int l(x,y)d\mu_{y}(y)$ (for any $\mu_{y}$) and $y\mapsto\int l(x,y)d\mu_{x}(x)$ (for any $\mu_{x}$). By the convergence $\gamma^{n}\to\gamma$, for a given $\tilde{\varepsilon}>0$, there is a $N_{\tilde{\varepsilon}}\in\mathbb{N}$ such that $d_{BL}(\gamma^{n},\gamma)\leq\tilde{\varepsilon}$ a.s. for all $n\geq N_{\tilde{\varepsilon}}$. This also implies the same property for the $t$-marginals: a.s., for any $t\in[0,T]$, $d_{BL}(\gamma^{n}_{t},\gamma_{t})\leq\tilde{\varepsilon}$ if $n\geq N_{\tilde{\varepsilon}}$. Therefore, since the constant $\tilde{C}$ does not depend on $n$ or $t$,